Fill in the blanks.
 (b) 13 + _____ = 0
(c) 12 + (–12) = ____
 
 
iaemfuewgnuersiojmhnbiorjhmneru

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student


(b) 13 + (-13) = 0

(c) 12 + (-12) = 0


​Thanks
nanditha jampala
2 years ago
Given Question:

 

Fill in the blanks.
 (b) 13 + _____ = 0
(c) 12 + (–12) = ____
 
 
Answer:
(a) 13 + (-13) = 0
(b) 12 + (-12) = 0
 
thank you